Calculate the number of milliliters of 0.611 M Ba(OH)2 required to precipitate all of the Ca2+ ions in 101 mL of 0.762 M CaCl2 solution as Ca(OH)2. The equation for the reaction is:


CaCl2(aq) + Ba(OH)2(aq) Ca(OH)2(s) + BaCl2(aq)


 mL Ba(OH)2




1
Expert's answer
2021-10-11T06:15:56-0400

Molar mass of Ba(OH)2= 171.34 g/mol

171.34×0.762= 130.56g

171.34×0.611= 104.69g

=130.36/104.69

=1.245ml